(a) The Ideal Gas Law equation is derived from the three gas laws. Explain the

THREE gas laws by providing mathematical equations showing the

relationship between the relevant quantities.


(b) A closed gas cylinder contains 0.500 mol H2, 0.300 mol O2, 1.200 mol CO2 at a

temperature of 25 °C and a pressure of 2.00 atm.

i. Explain Dalton’s Law of partial pressure.

ii. Calculate the volume of the cylinder.

iii. Calculate the partial pressure of O2 in the container


1
Expert's answer
2021-04-25T03:48:52-0400

a. The three gas laws are Boyle's law, Charles law and Avogadro's law.

Boyle's law states that the volume of a given mass of a gas is inversely proportional to its pressure at constant Temperature.

PV=K

Charles law states that the volume of a given mass of a gas is directly proportional to it's temperature at constant pressure.

V/T=K

Avogadro's law states that equal volumes of all gases at the same temperature and pressure contains the same number of moles.

n/V=K

Combining the three Equations give the ideal gas equation PV=nRT


b. I) Dalton's law of partial pressure states that the total pressure in a container containing a mixture of gases is equal to the sum of the partial pressures exerted by each gas.

II) Total mole= 0.500+0.300+1.200=2.000 mole

From ideal gas equation

PV=nRT

V=nRT/P

V= 2.000x0.0821x298/2.0

V=24.47L


III) Po= mole fraction of oxygen x total pressure

Po= 0.300/2.000 x 2.0 = 0.30atm